# Visibility into AI Agents

### v5

@@ A.1. The Underlying System
  • Evaluations of the system’s degree of agency [Mialon et al., 2023, Ruan et al., 2023, Liu et al., 2023];
  • Evaluations of the system’s generality: its ability to accomplish a broad array of tasks to some specified performance threshold [Morris et al., 2023];
  • Red flags, such as previous incidents or results of previous dangerous capabilities and alignment evaluations [Shevlane et al., 2023];
  • Dependencies of the agent, such as with an ecosystem graph [Bommasani et al., 2023] (e.g., whether the agent is a fine-tuned variant of another model).
@@ A.2. The Specific Instance of the Agent
  • The agent’s goal, including both what the user specifies and what the system appears to be achieving [Gilbert et al., 2023];

### v6

@@ A.1. The Underlying System
  • Evaluations of the system’s degree of agency (Mialon et al., 2023; Ruan et al., 2023; Liu et al., 2023);
  • Evaluations of the system’s generality: its ability to accomplish a broad array of tasks to some specified performance threshold (Morris et al., 2023);
  • Red flags, such as previous incidents or results of previous dangerous capabilities and alignment evaluations (Shevlane et al., 2023);
  • Dependencies of the agent, such as with an ecosystem graph (Bommasani et al., 2023) (e.g., whether the agent is a fine-tuned variant of another model).
@@ A.2. The Specific Instance of the Agent
  • The agent’s goal, including both what the user specifies and what the system appears to be achieving (Gilbert et al., 2023);